OpenStudy (anonymous):

f(x)=(x+5)^2(x-3)^2 for intercepts I got: x (-5,0),(3,0) y (0,225)

OpenStudy (callisto):

Hmm... May I know what your question is?

OpenStudy (anonymous):

did i do this correct

OpenStudy (callisto):

Yes. For x-int, put f(x) = 0 0 = (x+5)^2(x-3)^2 (x+5)^2=0 or (x-3)^2=0 x = -5 or x = 3 So, x-int are at ( -5, 0) and (3, 0) For y-int, y-int = f(0) = (0+5)^2(0-3)^2 = 5^2 ( -3)^2 = 25 (9) = 225 So, y-int is at (0, 225)

OpenStudy (anonymous):

thank you

OpenStudy (callisto):

Welcome :)